OpenStudy (anonymous):

Solve the equation. Show all work please. 9w-4=-6

jimthompson5910 (jim_thompson5910):

9w-4=-6 9w=-6+4 9w=-2 w=-2/9

OpenStudy (saifoo.khan):

\[w = -\frac29\]

OpenStudy (anonymous):

cheater

OpenStudy (saifoo.khan):

Proud to be the one.

OpenStudy (anonymous):

not you saifoo

OpenStudy (anonymous):

The other guy

OpenStudy (saifoo.khan):

I wasn't referring to u,.

OpenStudy (anonymous):

love yall:)